OpenStudy (anonymous):

The function f(x) = 245(4)^x represents the growth of a fruit fly population every year in the orange grove. Stacey wants to manipulate the formula to an equivalent form that calculates every month, not every year. Which function is correct for Stacey's purposes?

OpenStudy (jdoe0001):

well.. there are 12 months in 1 year so for any value, one twelth of that will be what every month averages to \(\bf f(x)=245(4)^x\impliedby yearly \\ \quad \\ f(x)=\cfrac{245(4)^x}{12}\impliedby monthly\)

OpenStudy (anonymous):

alright these are the answer choices f(x) = 45(4)^x f(x) = 45(4^12)^x/12 f(x) = 2752(4)x f(x) = 245(4 ^1/2)^12x

OpenStudy (anonymous):

@jdoe0001 based on ur response i'd say it's either B or D

OpenStudy (kropot72):

Please confirm that choice D should be written as: \[\large 245(4^{\frac{1}{12}})^{12x}\]

OpenStudy (anonymous):

@kropot72 yes it's written exactly like that

OpenStudy (kropot72):

Well, that is the correct choice.

OpenStudy (condor34):

kropot could you help me when you are done

OpenStudy (kropot72):

@Condor34 Sorry, maybe later. I must log out now.

OpenStudy (anonymous):

thank you!!! i was unsure because B & D seemed similar

OpenStudy (kropot72):

You're welcome :)